An object weighing 58,800 N rests on the flatbed of a truck. It is held in place by metal brackets that can exert a maximum horizontal force of 9000 N. When the truck is traveling 15 m/s, what is the minimum stopping distance if the load is not to slide forward into the cab?

1
Expert's answer
2022-03-15T10:36:32-0400

Solution.

"P=58800N;"

"F=9000N;"

"v0=15m\/s;"

"P=mg\implies m=\dfrac{P}{g};"

"m=\dfrac{58800}{9.8}=6000kg;"

"F=ma\implies a=\dfrac{F}{m};"

"a=\dfrac{9000}{6000}=1.5 m/s;"

"S=\dfrac{v^2-v_0^2}{2a};"



"S=\dfrac{-225}{-3}=75m;"

Answer: "S=75m."